# Characterizations of voting rules based on majority margins

### v2

@@ Appendix A Appendix: Allowing Noncomparability
  • if $n=\#P_{\mathbf{R}}(x,y)-\#P_{\mathbf{Q}}(x,y)>0$, then since $\mathcal{M}_{\mathbf{R}}=\mathcal{M}_{\mathbf{Q}}$, #$\#P_{\mathbf{R}}(y,x)-\#P_{\mathbf{Q}}(y,x)$ is also $n$, and in this case add to $\mathbf{Q}_{1}$ $n$ pairs of fresh voters where each pair votes $R_{x>y}$ and $R_{y>x}$;
  • if $n=\#P_{\mathbf{Q}}(x,y)-\#P_{\mathbf{R}}(x,y)>0$, then since $\mathcal{M}_{\mathbf{R}}=\mathcal{M}_{\mathbf{Q}}$, #$\#P_{\mathbf{Q}}(y,x)-\#P_{\mathbf{R}}(y,x)$ is also $n$, and in this case add to $\mathbf{R}_{1}$ $n$ pairs of fresh voters where each pair votes $R_{x>y}$ and $R_{y>x}$.
It is clear that $\mathbf{R}_{1}$ and $\mathbf{Q}_{1}$ have not only the same margins but also the same #$\#P$ function. They may have different numbers of voters, but we can also add to the smaller profile enough fresh voters with ballots having no strict preferences (i.e., the identity relation) and obtain $\mathbf{R}_{2}$ and $\mathbf{Q}_{2}$. Then $\mathcal{H}(\mathbf{R}_{2})=\mathcal{H}(\mathbf{Q}_{2})$. Using the assumption that $F$ is head-to-head and satisfies Nonlinear Neutral Reversal and Neutral Blankness, we have $F(\mathbf{R})=F(\mathbf{Q})$.
For the case where $D=\mathsf{WOSN}$, again we take profiles $\mathbf{R}$ and $\mathbf{Q}$ from $\mathsf{WOSN}$ such that $\mathcal{M}_{\mathbf{R}}=\mathcal{M}_{\mathbf{Q}}$, and we want to show that $F(\mathbf{R})=F(\mathbf{Q})$. As in the above case, we construct $\mathbf{R}_{1}$ and $\mathbf{Q}_{1}$ in the same way. Now we also need to modify their #$\#I$ functions to be the same. This can be done by the self-reversing WOSN ballot $R_{x=y}$ where there is no strict preference and the only non-trivial indifference is between $x$ and $y$. In other words, $R_{x=y}$ is the union of the identity relation on $X(\mathbf{R})$ and $\{(x,y),(y,x)\}$. We obtain $\mathbf{R}_{2}$ and $\mathbf{Q}_{2}$ by adding, for each $x,y\in X$ with $x\prec y$, $n=\#I_{\mathbf{R}_{1}}(x,y)-\#I_{\mathbf{Q}_{1}}(x,y)$ many fresh $L_{x=y}$ ballots to $\mathbf{Q}_{1}$ if $n$ is positive, or $m=\#I_{\mathbf{Q}_{1}}(x,y)-\#I_{\mathbf{R}_{1}}(x,y)$ many fresh $L_{x=y}$ ballots to $\mathbf{R}_{1}$ if $m$ is positive. Then both the #$\#P$ function and the #$\#I$ function of $\mathbf{R}_{2}$ and $\mathbf{Q}_{2}$ are the same. To make the #$\#N$ functions the same as well, we only need to add blank ballots to the smaller profile and obtain $\mathbf{R}_{3}$ and $\mathbf{Q}_{3}$ of equal size. Since $F$ is head-to-head and $\mathcal{H}(\mathbf{R}_{3})=\mathcal{H}(\mathbf{Q}_{3})$, we have $F(\mathbf{R}_{3})=F(\mathbf{Q}_{3})$. Then since $F$ satisfies Neutral Self-Reversal and Nonlinear Neutral Reversal, $F(\mathbf{R})=F(\mathbf{Q})$. ∎

### v3

@@ Appendix A Appendix: Allowing Noncomparability
  • if $n=\#P_{\mathbf{R}}(x,y)-\#P_{\mathbf{Q}}(x,y)>0$, then since $\mathcal{M}_{\mathbf{R}}=\mathcal{M}_{\mathbf{Q}}$, $P_{\mathbf{R}}(y,x)-\#P_{\mathbf{Q}}(y,x)$ is also $n$, and in this case add to $\mathbf{Q}_{1}$ $n$ pairs of fresh voters where each pair votes $R_{x>y}$ and $R_{y>x}$;
  • if $n=\#P_{\mathbf{Q}}(x,y)-\#P_{\mathbf{R}}(x,y)>0$, then since $\mathcal{M}_{\mathbf{R}}=\mathcal{M}_{\mathbf{Q}}$, $P_{\mathbf{Q}}(y,x)-\#P_{\mathbf{R}}(y,x)$ is also $n$, and in this case add to $\mathbf{R}_{1}$ $n$ pairs of fresh voters where each pair votes $R_{x>y}$ and $R_{y>x}$.
It is clear that $\mathbf{R}_{1}$ and $\mathbf{Q}_{1}$ have not only the same margins but also the same #P\#P function. They may have different numbers of voters, but we can also add to the smaller profile enough fresh voters with ballots having no strict preferences (i.e., the identity relation) and obtain $\mathbf{R}_{2}$ and $\mathbf{Q}_{2}$. Then $\mathcal{H}(\mathbf{R}_{2})=\mathcal{H}(\mathbf{Q}_{2})$. Using the assumption that $F$ is head-to-head and satisfies Nonlinear Neutral Reversal and Neutral Blankness, we have $F(\mathbf{R})=F(\mathbf{Q})$.
For the case where $D=\mathsf{WOSN}$, again we take profiles $\mathbf{R}$ and $\mathbf{Q}$ from $\mathsf{WOSN}$ such that $\mathcal{M}_{\mathbf{R}}=\mathcal{M}_{\mathbf{Q}}$, and we want to show that $F(\mathbf{R})=F(\mathbf{Q})$. As in the above case, we construct $\mathbf{R}_{1}$ and $\mathbf{Q}_{1}$ in the same way. Now we also need to modify their #I\#I functions to be the same. This can be done by the self-reversing WOSN ballot $R_{x=y}$ where there is no strict preference and the only non-trivial indifference is between $x$ and $y$. In other words, $R_{x=y}$ is the union of the identity relation on $X(\mathbf{R})$ and $\{(x,y),(y,x)\}$. We obtain $\mathbf{R}_{2}$ and $\mathbf{Q}_{2}$ by adding, for each $x,y\in X$ with $x\prec y$, $n=\#I_{\mathbf{R}_{1}}(x,y)-\#I_{\mathbf{Q}_{1}}(x,y)$ many fresh $R_{x=y}$ ballots to $\mathbf{Q}_{1}$ if $n$ is positive, or $m=\#I_{\mathbf{Q}_{1}}(x,y)-\#I_{\mathbf{R}_{1}}(x,y)$ many fresh $R_{x=y}$ ballots to $\mathbf{R}_{1}$ if $m$ is positive. Then both the #P\#P function and the #I\#I function of $\mathbf{R}_{2}$ and $\mathbf{Q}_{2}$ are the same. To make the #N\#N functions the same as well, we only need to add blank ballots to the smaller profile and obtain $\mathbf{R}_{3}$ and $\mathbf{Q}_{3}$ of equal size. Since $F$ is head-to-head and $\mathcal{H}(\mathbf{R}_{3})=\mathcal{H}(\mathbf{Q}_{3})$, we have $F(\mathbf{R}_{3})=F(\mathbf{Q}_{3})$. Then since $F$ satisfies Neutral Self-Reversal and Nonlinear Neutral Reversal, $F(\mathbf{R})=F(\mathbf{Q})$. ∎